Match the following aqueous solutions with the appropriate letter from the column on the right.

How does the m reflect boiling point?

1. 0.14 m CrI2    A. Highest boiling point
2. 0.13 m Na2CO3 B. Second highest boiling point
3. 0.15 m ZnCl2 C. Third highest boiling point
4. 0.47 m Ethylene glycol(nonelectrolyte) D. Lowest boiling point
Answer #1

There will an increase in boiling point of water when solute
is added to it.it is given by
dtb = i.Kb.m
Kb= constant for water, i= factor depends on number of ions,m=molality


1.   0.14 m   CrI2
here m= 0.14, i= 3 [3 ions forms]
dTb = 3*0.14 *Kb = 0.42 Kb
      
2.   0.13 m   Na2CO3
m= 0.13, i= 3 [3 ions forms]  
dTb= 0.13*3.Kb = 0.39 Kb  

3.   0.15 m   ZnCl2
m= 0.15 , i= 3
dTb = 0.45Kb  

  
4.   0.47 m   Ethylene glycol

m=0.47 , i=1 [it will ot ionize]
dTb = 0.47 Kb **

So highest dTb value means highest boling point . Hence glycol
has highest BP

4.0.47m Ethylene glycol.. Highest boiling point

3.0.15 m ZnCl2...Second highest boiling point

1.0.14 m CrI2....Third highest boiling point

2. 0.13m Na2CO3.....Lowest boiling point